Two portfolios, P1 and P2, produce the following returns in years 1-5:

Year P1 Return (%) P2 Return (%)
1 8 -10
2 3 -4
3 -6 4
4 7 -6
5 7 5

What is the mean annual return on a portfolio that is 70% invested in P1 and 30% invested in P2?

Enter answer accurate to 2 decimal places
Answer #1

Average return on P1 = (8+3-6+7+7)/ 5 = 3.8%

Average return on P2 = (-10-4+4-6+5)/5 = -2.2%

Mean return on a portfolio = Weight of P1* Average return on P1 + weight of P2*Average return on P2

= 0.7*3.8%+ 0.3*-2.2%

=2%